# Improved multifidelity Monte Carlo estimators based on normalizing flows and dimensionality reduction techniques

### v1

@@ Abstract
We study the problem of multifidelity uncertainty propagation for computationally expensive models. In particular, we consider the general setting where the high-fidelity and low-fidelity models have a dissimilar parameterization both in terms of number of random inputs and their probability distributions, which can be either known in closed form or provided through samples. We derive novel multifidelity Monte Carlo estimators which rely on a shared subspace between the high-fidelity and low-fidelity models where the parameters follow the same probability distribution, i.e., a standard Gaussian. We build the shared space employing normalizing flows to map different probability distributions into a common one, together with linear and nonlinear dimensionality reduction techniques, active subspaces and autoencoders, respectively, which capture the subspaces where the models vary the most. We then compose the existing low-fidelity model with these transformations and construct modified models with an increased correlation with the high-fidelity, which therefore yield multifidelity Monte Carlo estimators with reduced variance. A series of numerical experiments illustrate the properties and advantages of our approaches.
@@ 1 Introduction
Uncertainty quantification has become a crucial component of computational modeling, as a way to enhance the validity and utility of numerical simulations. Uncertainty quantification studies can provide confidence metrics for quantities of interest and inform future data collection through sensitivity and identifiability analysis. However, in many cases, a naive approach to uncertainty quantification quickly becomes computationally infeasible as a result of the large computational cost needed to numerically solve complex physics-based mathematical models. Therefore, maintaining a reasonable computational cost for an uncertainty quantification study becomes challenging when relying solely on high-fidelity simulations. This has motivated the development of multilevel and multifidelity Monte Carlo strategies that offset the computational cost of estimation to low-fidelity models, accelerating convergence and improving the tractability of uncertainty quantification for computationally expensive simulations [11, 23, 25, 12].
@@
In some cases, however, the low fidelity-model can be obtained through a substantial simplification of the high-fidelity model, resulting in input parameters with potentially both different dimensionality and probability distribution, which we refer to as dissimilar parameterization. We remark that in this case the performance of standard multifidelity Monte Carlo estimators decreases, and it is important to concentrate the variability in few dimensions to enhance the correlation [10, 39]. We therefore propose two methodologies to create a shared subspace of reduced dimension which acts as a bridge between the two models. The first step in obtaining the shared subspace consists of finding an “important” subspace individually for both the high-fidelity and low-fidelity models where they vary the most, and consequently which captures most of the variance of the models. Perturbations of the random inputs along such important directions are responsible for most of the variability in the model response. If such a structure is present in the problem, then we can consider model responses only for inputs in the subspace, effectively reducing the problem dimensionality. In this work we employ two different dimensionality reduction techniques, active subspaces and autoencoders, which provide linear and nonlinear transformations, respectively. Active subspaces have been first formalized in [3, 4], while an introduction about autoencoder for unsupervised learning can be found in [2]. Autoencoders have become popular in the past few years thanks to their expressibility, as they leverage neural networks to find nonlinear transformations of the data, as opposed to linear maps provided by active subspaces. The second ingredient to obtain a shared subspace is a normalizing flow, i.e., an invertible transformation from a generic probability distribution into an easy-to-sample base distribution, usually a standard Gaussian. For a comprehensive review about normalizing flows we refer to [19, 24]. The goal of normalizing flows is enforcing the same probability distribution, in particular a standard Gaussian, of the latent variables, i.e., the parameters in the shared space. For active subspaces, which are known to preserve Gaussianity, we build normalizing flows from the input distributions of the parameters, such that the reduced subspaces of both the high-fidelity and low-fidelity models are automatically Gaussian. On the other hand, for autoencoders we learn normalizing flows from the distributions of the latent variables of the model into the shared subspace, which is therefore Gaussian. We remark that these subspaces and normalizing flows are approximated from the results of a pilot run that is typically employed as a first step in multifidelity Monte Carlo estimation, without the need to run additional high-fidelity simulations.
@@
We reasonably assume that if the variability of the high-fidelity and low-fidelity models is concentrated in few variables, then using the shared space as a bridge aligning the important directions of the fidelities would improve the correlations and consequently the multifidelity estimator. Starting from the existing low-fidelity model, we therefore construct new low-fidelity models whose inputs are the same parameters of the high-fidelity model which are transformed into inputs for the original low-fidelity model in such a way that the correlation between the fidelities increases. We finally obtain multifidelity estimators which are unbiased, as the high-fidelity model does not change, and with reduced variance with respect to standard multifidelity Monte Carlo estimators. The present work generalizes and extends a series of contributions in this area like, e.g., [9, 10, 39], where the challenge of dissimilar parameterization has been tackled with either active subspaces or adaptive basis [36]. The main extension regarding previous work on active subspaces is that we include a normalizing flow which allows for any input distribution of the parameters, even known only through samples. Moreover, both active subspaces and adaptive basis are linear dimension reduction strategies, and therefore autoencoders are a natural extension since they provide nonlinear transformations.
@@ 2 Review of current methods
In this section we briefly review the main tools employed in our uncertainty quantification pipelines. We recall that our goal is the efficient estimation of scalar quantities of interest of computationally expensive models. Let $Q\colon\mathbb{R}^{d}\to\mathbb{R}$ represent a computational model, and let $\bm{\xi}\in\mathbb{R}^{d}$ be a random vector of inputs distributed according to the joint distribution $\mu$. We aim to characterize the statistical moments of the quantity of interest $Q(\bm{\xi})$, focusing in particular on its expectation $\operatorname{\mathbb{E}}[Q(\bm{\xi})]$. The standard Monte Carlo estimator approximates this expected value through a set of $N$ realizations $\{\bm{\xi}_{n}\}_{n=1}^{N}$ of the input variable $\bm{\xi}\sim\mu$ as
@@ 2.1 Multifidelity Monte Carlo estimator
Let us assume that a computationally cheap model for the original model $Q$ is available, and denote $Q^{\mathrm{HF}}$ and $Q^{\mathrm{LF}}$ the original (high-fidelity) and the cheap (low-fidelity) models, respectively. We notice that the low-fidelity model can be any, even biased, approximation of the high-fidelity model, as long as it is computationally cheap to simulate. We adopt the estimator originally introduced in [23], which focus on the case of a single low-fidelity model, but the multifidelity Monte Carlo estimator can be easily extended, or generalized, to the case of multiple low-fidelity models, e.g.,  [25, 12]. Let $w=\mathcal{C}^{\mathrm{LF}}/\mathcal{C}^{\mathrm{HF}}$ be the cost ratio between the two fidelities, and let $\mathcal{B}$ be the computational budget available in terms of high-fidelity evaluations, i.e.,
@@
$\operatorname{\mathbb{V}ar}\left[\widehat{Q}_{N^{\mathrm{HF}},N^{\mathrm{LF}}}^{\mathrm{MFMC}}\right]=\operatorname{\mathbb{V}ar}\left[\widehat{Q}_{\mathcal{B}}^{\mathrm{HF},\mathrm{MC}}\right]\left(\sqrt{1-\rho^{2}}+\sqrt{w\rho^{2}}\right),$ (2.8)
@@ 2.2 Active subspaces
Note that since $\widetilde{C}$ is a symmetric positive semidefinite matrix, then it admits a real eigenvalue decomposition
@@
where $W$ is orthogonal and its columns are the eigenvectors of $C$, and $\Lambda$ is a diagonal matrix which contains the corresponding eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{d}\geq 0$, which can be arranged in decreasing order. This ordering suggests a possible separation between important (or active) and irrelevant (or inactive) contribution of the corresponding eigenvector to the linear decomposition of $C$ as a sum of rank-one matrices, in particular if the smallest eigenvalues are close to zero. To better visualize such separation, we write
@@
where $\Lambda_{A}\in\mathbb{R}^{r\times r}$ and $W_{A}\in\mathbb{R}^{d\times r}$ contain the first $r<d$ eigenvalues and eigenvectors. The column spans of $W_{A}$ and $W_{I}$ represent the active and inactive subspace, respectively. These linear transformations allow us to decompose the original input $\bm{\xi}\in\mathbb{R}^{d}$ into the active and inactive parts as follows
@@
This suggests the possibility to generate shared parameterizations even for models having a different number of random inputs, as long as their active inputs share the same dimensionality and distribution [9, 10]. For general probability distributions, it is often possible to define a transformation mapping $\mu$ to a standard Gaussian via normalizing flows, compute the active subspace, and finally transform the variables back to their original distribution after sampling. We notice that this is a significant advantage since it allows one to handle any type of input data, independently of their correlation. As observed in [9], even if a transformation introduces additional complexity, the increase in computational cost could be outweighed by the increase in correlation between models. A strong limitation of the active subspace technique is that it provides only linear maps for dimensionality reduction. This restriction can be overcome by replacing active subspaces with autoencoders, which allow for nonlinear transformations. The next two sections will therefore focus on autoencoders and normalizing flows, respectively.
@@ 2.3 Autoencoders
Autoencoders are a data-driven approach widely used for unsupervised dimensionality reduction, with the ability to learn an intrinsic structure existing in the data by leveraging neural networks. They consist of one encoder $\mathcal{E}$ followed by one decoder $\mathcal{D}$, where the encoder $\mathcal{E}\colon\mathbb{R}^{d}\to\mathbb{R}^{r}$ with $r<d$ compresses the original input $\bm{\xi}\in\mathbb{R}^{d}$ into a latent representation $\bm{x}\in\mathbb{R}^{r}$, and the decoder $\mathcal{D}\colon\mathbb{R}^{r}\to\mathbb{R}^{d}$ seeks to reconstruct the original input $\bm{\xi}$ starting from the latent representation $\bm{x}$, in the sense that $\mathcal{D}(\mathcal{E}(\bm{\xi}))\simeq\bm{\xi}$. The functions $\mathcal{E}$ and $\mathcal{D}$ are usually parameterized by fully connected neural networks. In this work we are not interested in reconstructing exactly the original input, but rather $Q(\bm{\xi})$. In particular, given a model $Q$, analogously to the active subspace approach, we would like to find a lower dimensional representation of $Q$ by selecting a manifold of dimension $r$ where the function varies the most. Hence, we aim to reconstruct the original quantity of interest, meaning that $Q(\mathcal{D}(\mathcal{E}(\bm{\xi})))\simeq Q(\bm{\xi})$. We notice that this can be seen as a supervised dimensionality reduction, where we would like to learn a structure in the data according to some metric, which in this case is the model $Q$. We parameterize the encoder and the decoder as fully connected neural networks $\mathcal{E}(\cdot;\phi_{E})$ and $\mathcal{D}(\cdot;\phi_{D})$ with hyperbolic tangent activation function, and we compute the optimal parameters by minimizing the loss function
@@
where $\{\bm{\xi}_{n}\}_{n=1}^{N}$ is the sample of available data. We notice that minimizing the loss function requires multiple evaluations of the model $Q$, and this can be impractical if the model is computationally expensive. We therefore build a surrogate model which is only used in the training process, as we discuss in Remark 3.3. Using the same terminology as for the active subspaces technique, we say that the active variable is $\bm{\xi}_{A}=\mathcal{E}(\bm{\xi})$ and we notice that its distribution is unknown. Therefore, by constructing a map from the probability distribution of the active variable into a standard Gaussian through normalizing flows (see Section 2.4), we can generate a shared parameterization even for models having a different number of inputs, as long as the reduced dimension of the autoencoder is the same. We only remark that, since the autoencoders do not have a ranking of the latent dimension like the active subspace technique, if the latent space is multidimensional, then we also have to find the best permutation in term of correlation of the components of the latent variable.
@@ 2.4 Normalizing flows
Normalizing flows are invertible transformations which map generic probability distributions into more simple and tractable distributions, usually standard Gaussian. Let $\mu$ be a probability distribution on $\mathbb{R}^{d}$ with density $\psi$ and let $\mu_{0}$ with density $\psi_{0}$ be the target distribution, which in our work, like in most of the cases, is $\mu_{0}=\mathcal{N}(\bm{0},I_{d})$. We aim to find a diffeomorphism $\mathcal{T}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $\mathcal{T}_{\#}\mu=\mu_{0}$, where $\mathcal{T}_{\#}$ denotes the pushforward measure through the map $\mathcal{T}$. We consider a parameterization $\mathcal{T}(\cdot;\theta)$, which is the composition of invertible transformations defined by neural networks. Then, given a sample $\{\bm{\xi}_{n}\}_{n=1}^{N}$ from the initial distribution $\mu$, the best parameters $\theta$ are computed by maximizing the log-likelihood function for the density $\widetilde{\psi}(\cdot;\theta)$ of the measure $\mathcal{T}^{-1}(\cdot;\theta)_{\#}\mu_{0}$. In practice, we minimize the loss function

### v2

@@ Abstract
We study the problem of multifidelity uncertainty propagation for computationally expensive models. In particular, we consider the general setting where the high-fidelity and low-fidelity models have a dissimilar parameterization both in terms of number of random inputs and their probability distributions, which can be either known in closed form or provided through samples. We derive novel multifidelity Monte Carlo estimators which rely on a shared subspace between the high-fidelity and low-fidelity models where the parameters follow the same probability distribution, i.e., a standard Gaussian. We build the shared space employing normalizing flows to map different probability distributions into a common one, together with linear and nonlinear dimensionality reduction techniques, active subspaces and autoencoders, respectively, which capture the subspaces where the models vary the most. We then compose the existing low-fidelity model with these transformations and construct modified models with an increased correlation with the high-fidelity model, which therefore yield multifidelity estimators with reduced variance. A series of numerical experiments illustrate the properties and advantages of our approaches.
@@ 1 Introduction
Uncertainty quantification has become a crucial component of computational modeling, as a way to enhance the validity and utility of numerical simulations. Uncertainty quantification studies can provide confidence metrics for quantities of interest and inform future data collection through sensitivity and identifiability analysis. However, in many cases, a naive approach to uncertainty quantification quickly becomes computationally infeasible as a result of the large computational cost needed to numerically solve complex physics-based mathematical models. Therefore, maintaining a reasonable computational cost for an uncertainty quantification study becomes challenging when relying solely on high-fidelity simulations. This has motivated the development of multilevel and multifidelity Monte Carlo strategies that offset the computational cost of estimation to low-fidelity models, accelerating convergence and improving the tractability of uncertainty quantification for computationally expensive simulations [14, 27, 29, 15, 3, 34, 35, 6].
@@
In some cases, however, the low fidelity-model can be obtained through a substantial simplification of the high-fidelity model, resulting in input parameters with potentially both different dimensionality and probability distribution, which we refer to as dissimilar parameterization. We remark that in this case the performance of standard multifidelity Monte Carlo estimators decreases, and it is important to concentrate the variability in few dimensions to enhance the correlation [12, 45]. We therefore propose two methodologies to create a shared subspace of reduced dimension which acts as a bridge between the two models. The first step in obtaining the shared subspace consists of finding an “important” subspace individually for both the high-fidelity and low-fidelity models where they vary the most, and consequently which captures most of the variance of the models. Perturbations of the random inputs along such important directions are responsible for most of the variability in the model response. If such a structure is present in the problem, then we can consider model responses only for inputs in the subspace, effectively reducing the problem dimensionality. In this work we employ two different dimensionality reduction techniques, active subspaces and autoencoders, which provide linear and nonlinear transformations, respectively. Active subspaces have been first formalized in [4, 5], while an introduction about autoencoder for unsupervised learning can be found in [2]. Autoencoders have become popular in the past few years thanks to their expressibility, as they leverage neural networks to find nonlinear transformations of the data, as opposed to linear maps provided by active subspaces. The second ingredient to obtain a shared subspace is a normalizing flow, i.e., an invertible transformation from a generic probability distribution into an easy-to-sample base distribution, usually a standard Gaussian. For a comprehensive review about normalizing flows we refer to [22, 28]. The goal of normalizing flows is enforcing the same probability distribution, in particular a standard Gaussian, of the latent variables, i.e., the parameters in the shared space. For active subspaces, which are known to preserve Gaussianity, we build normalizing flows from the input distributions of the parameters, such that the reduced subspaces of both the high-fidelity and low-fidelity models are automatically Gaussian. On the other hand, for autoencoders we learn normalizing flows from the distributions of the latent variables of the model into the shared subspace, which is therefore Gaussian. We remark that these subspaces and normalizing flows are approximated from the results of a pilot run that is typically employed as a first step in any multifidelity estimator, without the need to run additional high-fidelity simulations.
@@
We reasonably assume that if the variability of the high-fidelity and low-fidelity models is concentrated in few variables, then using the shared space as a bridge aligning the important directions of the fidelities would improve the correlations and consequently reduce the variance of the multifidelity estimator. Starting from the existing low-fidelity model, we therefore construct new low-fidelity models whose inputs are the same parameters of the high-fidelity model which are transformed into inputs for the original low-fidelity model in such a way that the correlation between the fidelities increases. We finally obtain multifidelity estimators which are unbiased, as the high-fidelity model does not change, and with reduced variance with respect to standard multifidelity Monte Carlo estimators. The present work generalizes and extends a series of contributions in this area like, e.g., [11, 12, 45], where the challenge of dissimilar parameterization has been tackled with either active subspaces or adaptive basis [42]. The main extension regarding previous work on active subspaces is that we include a normalizing flow which allows for any input distribution of the parameters, even known only through samples. Moreover, both active subspaces and adaptive basis are linear dimension reduction strategies, and therefore autoencoders are a natural extension since they provide nonlinear transformations.
@@ 2 Review of current methods
In this section we briefly review the main tools employed in our uncertainty quantification pipelines. We recall that our goal is the efficient estimation of scalar quantities of interest of computationally expensive models. Let $Q\colon\mathbb{R}^{d}\to\mathbb{R}$ represent a computational model, and let $\bm{\xi}\in\mathbb{R}^{d}$ be a random vector of inputs distributed according to the joint distribution $\mu$. We aim to characterize the statistical moments of the quantity of interest $Q(\bm{\xi})$, focusing in particular on its expectation $\operatorname{\mathbb{E}}[Q(\bm{\xi})]$. For higher-order moments, we can just replace the quantity of interest $Q(\bm{\xi})$ by its power $Q^{m}(\bm{\xi})$ with $m>1$. The standard Monte Carlo estimator approximates this expected value through a set of $N$ realizations $\{\bm{\xi}_{n}\}_{n=1}^{N}$ of the input variable $\bm{\xi}\sim\mu$ as
@@ 2.1 Multifidelity Monte Carlo estimator
Let us assume that a computationally cheap model for the original model $Q$ is available, and denote $Q^{\mathrm{HF}}$ and $Q^{\mathrm{LF}}$ the original (high-fidelity) and the cheap (low-fidelity) models, respectively. We notice that the low-fidelity model can be any, even biased, approximation of the high-fidelity model, as long as it is computationally cheap to simulate. We adopt the estimator originally introduced in [27], which focus on the case of a single low-fidelity model, but the multifidelity Monte Carlo estimator can be easily extended, or generalized, to the case of multiple low-fidelity models, e.g.,  [29, 15, 3, 34, 35, 6]. Let $w=\mathcal{C}^{\mathrm{LF}}/\mathcal{C}^{\mathrm{HF}}$ be the cost ratio between the two fidelities, and let $\mathcal{B}$ be the computational budget available in terms of high-fidelity evaluations, i.e.,
@@
$\operatorname{\mathbb{V}ar}\left[\widehat{Q}_{N^{\mathrm{HF}},N^{\mathrm{LF}}}^{\mathrm{MFMC}}\right]=\operatorname{\mathbb{V}ar}\left[\widehat{Q}_{\mathcal{B}}^{\mathrm{HF},\mathrm{MC}}\right]\left(\sqrt{1-\rho^{2}}+\sqrt{w\rho^{2}}\right)^{2},$ (2.8)
@@ 2.2 Active subspaces
We remark that the gradient of the model $Q$ is required in order to compute the matrix $C$. It can be approximated through finite differences, linear approximations, surrogate models, or any other technique to compute derivatives numerically. Additional details about how we deal with the gradient in our work are given in Remark 3.1. Moreover, we note that more sophisticated strategies could be introduced to reduce the data requirment of this step, especially for the high-fidelity model. For instance, the approximation of $C$ could be obtained via a multifidelity approach as demonstrated in [23]. Note that since $\widetilde{C}$ is a symmetric positive semidefinite matrix, then it admits a real eigenvalue decomposition
@@
where $W$ is orthogonal and its columns are the eigenvectors of $C$, and $\Lambda$ is a diagonal matrix which contains the corresponding eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{d}\geq 0$, which can be arranged in decreasing order. This ordering suggests a possible separation between important (or active) and irrelevant (or inactive) contributions of the corresponding eigenvector to the linear decomposition of $C$ as a sum of rank-one matrices, in particular if the smallest eigenvalues are close to zero. To better visualize such separation, we write
@@
where $\Lambda_{A}\in\mathbb{R}^{r\times r}$ and $W_{A}\in\mathbb{R}^{d\times r}$ contain the first $r<d$ eigenvalues and eigenvectors, respectively. The column spans of $W_{A}$ and $W_{I}$ represent the active and inactive subspace, respectively. These linear transformations allow us to decompose the original input $\bm{\xi}\in\mathbb{R}^{d}$ into the active and inactive parts as follows
@@
This suggests the possibility of generating shared parameterizations even for models having a different number of random inputs, as long as their active inputs share the same dimensionality and distribution [11, 12]. For general probability distributions, it is often possible to define a transformation mapping $\mu$ to a standard Gaussian via normalizing flows, compute the active subspace, and finally transform the variables back to their original distribution after sampling. We notice that this is a significant advantage since it allows one to handle any type of input data, independently of their correlation. As observed in [11], even if a transformation introduces additional complexity, the increase in computational cost could be outweighed by the increase in correlation between models. A strong limitation of the active subspace technique is that it provides only linear maps for dimensionality reduction. This restriction can be overcome by replacing active subspaces with autoencoders, which allow for nonlinear transformations. The next two sections will therefore focus on autoencoders and normalizing flows, respectively.
@@ 2.3 Autoencoders
Autoencoders are a data-driven approach widely used for unsupervised dimensionality reduction, with the ability to learn an intrinsic structure existing in the data by leveraging neural networks. They consist of one encoder $\mathcal{E}$ followed by one decoder $\mathcal{D}$, where the encoder $\mathcal{E}\colon\mathbb{R}^{d}\to\mathbb{R}^{r}$ with $r<d$ compresses the original input $\bm{\xi}\in\mathbb{R}^{d}$ into a latent representation $\bm{x}\in\mathbb{R}^{r}$, and the decoder $\mathcal{D}\colon\mathbb{R}^{r}\to\mathbb{R}^{d}$ seeks to reconstruct the original input $\bm{\xi}$ starting from the latent representation $\bm{x}$, in the sense that $\mathcal{D}(\mathcal{E}(\bm{\xi}))\simeq\bm{\xi}$. The functions $\mathcal{E}$ and $\mathcal{D}$ are usually parameterized by fully connected neural networks. In this work we are not interested in reconstructing exactly the original input, but rather $Q(\bm{\xi})$. In particular, given a model $Q$, analogously to the active subspace approach, we would like to find a lower dimensional representation of $Q$ by selecting a manifold of dimension $r$ where the function varies the most. Hence, we aim to reconstruct the original quantity of interest, meaning that $Q(\mathcal{D}(\mathcal{E}(\bm{\xi})))\simeq Q(\bm{\xi})$. We notice that this can be seen as a supervised dimensionality reduction, where we would like to learn a structure in the data according to some metric, which in this case is the model $Q$. We parameterize the encoder and the decoder as fully connected neural networks $\mathcal{E}(\cdot;\phi_{E})$ and $\mathcal{D}(\cdot;\phi_{D})$ with hyperbolic tangent activation functions, and we compute the optimal parameters by minimizing the loss function
@@
where $\{\bm{\xi}_{n}\}_{n=1}^{N}$ is the sample of available data. We notice that minimizing the loss function requires multiple evaluations of the model $Q$, and this can be impractical if the model is computationally expensive. We therefore build a surrogate model which is only used in the training process, as we discuss in Remark 3.3. Using the same terminology as for the active subspaces technique, we say that the active variable is $\bm{\xi}_{A}=\mathcal{E}(\bm{\xi})$ and we notice that its distribution is unknown. Therefore, by constructing a map from the probability distribution of the active variable into a standard Gaussian through normalizing flows (see Section 2.4), we can generate a shared parameterization even for models having a different number of inputs, as long as the reduced dimension of the autoencoder is the same.
@@ 2.4 Normalizing flows
Normalizing flows are invertible transformations which map generic probability distributions into more simple and tractable distributions, usually standard Gaussian. Let $\mu$ be a probability distribution on $\mathbb{R}^{d}$ with density $\psi$ and let $\mu_{0}$ with density $\psi_{0}$ be the target distribution, which in our work, like in most cases, is $\mu_{0}=\mathcal{N}(\bm{0},I_{d})$. We aim to find a diffeomorphism $\mathcal{T}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $\mathcal{T}_{\#}\mu=\mu_{0}$, where $\mathcal{T}_{\#}$ denotes the pushforward measure through the map $\mathcal{T}$. We consider a parameterization $\mathcal{T}(\cdot;\theta)$, which is the composition of invertible transformations defined by neural networks. Then, given a sample $\{\bm{\xi}_{n}\}_{n=1}^{N}$ from the initial distribution $\mu$, the best parameters $\theta$ are computed by maximizing the log-likelihood function for the density $\widetilde{\psi}(\cdot;\theta)$ of the measure $\mathcal{T}^{-1}(\cdot;\theta)_{\#}\mu_{0}$. In practice, we minimize the loss function